OpenStudy (anonymous):

Use Lagrange’s method to find the maximum and minimum values of f(x,y)=y^2-x^2 on the ellipse: (((x^2)/4)+(y^2)=1)

OpenStudy (anonymous):

Solution for the variable y= y=-1/2sq root 4-x^2 y= sqroot 4-x^2/2

OpenStudy (anonymous):

The alternate forms are x^2+4y^2=4 x^2+y^2-1=0 1/4(x^2+4y^2)=1